OpenStudy (anonymous):

Solve the equation 4x^2=36

OpenStudy (saifoo.khan):

\[x^2 = \frac{36}{4}\] \[x^2 = 9\]

OpenStudy (anonymous):

x=3, x=-3

OpenStudy (hoblos):

4x^2=36 x^2 = 36/4 =9 x=±3